OpenStudy (anonymous):

Hello, it's been a long time since I've been on here. Wonder is many things have changed since. Can anyone please help me with finding all soultions in the interval [0,2π ) to the trig equations? 1) 2 cos x=√3 Thank you.

OpenStudy (anonymous):

start with \[\cos(x)=\frac{\sqrt{3}}{2}\] and then see if you can remember which numbers work

OpenStudy (anonymous):

\[\frac{\pi}{6}\] works because \[\cos(\frac{\pi}{6})=\frac{\sqrt{3}}{2}\] go straight down on the unit circle and see that \[\frac{11\pi}{6}\] also works

OpenStudy (anonymous):

I have π /6, 11/π 6...i wonder if this is correct???

OpenStudy (anonymous):

Thank you, Satellite73! It's just that we are required to write everything out so I was just a bit thrown off with what I got. So I'm hoping that I'm also correct on my end with the π/6, 11/π 6. Thanks so much for your help! Have a great day!

OpenStudy (anonymous):

yw